The product of 5 times a number plus the product of 3 times the same number is 16.
Answer by waynest(281)

number = x
first number = 5x
second number = 3x
5x + 3x = 16
8x = 16
8x/8 = 16/8
x = 2
check:
5(2) + 3(2) = 16
10 + 6 = 16
16 = 16